# Angular bispectrum of matter number counts in cosmic structures

### v2

@@ Abstract
The bispectrum of galaxy number counts is a key probe of large-scale structure, offering insights into the initial conditions of the Universe, the nature of gravity, and cosmological parameters. In this work, we present the first full-sky computation of the angular bispectrum in second-order perturbation theory without invoking the Limber approximation, and formulated for finite redshift bins via window functions. To our knowledge, even the Newtonian part within this setup is novel. Building on this, we also include, up to second order in perturbation theory, the dynamical general relativistic and radiation effects, together with the leading relativistic projection effects. For simplicity, we neglect tracer bias and line-of-sight integrated contributions, however note that in particular the former can be straightforwardly incorporated within our framework. We evaluate the bispectrum contributions for two redshift bins, $1.75\leq z\leq 2.25$ and $0.55\leq z\leq 0.65$, and compare our theoretical prediction against relativistic light-cone simulations, with line-of-sight integral effects removed so as to enable direct consistency checks. As expected, we find that the Newtonian contributions are typically one or more orders of magnitudes larger than the relativistic signal across the entire spectrum for both redshifts. At $z=2$, we find that projection and dynamical relativistic effects have comparable amplitudes on large scales; somewhat unexpectedly, however, radiation effects dominate the relativistic signal in the squeezed limit. At $z=0.6$, the expected hierarchy is recovered, though dynamical corrections (from nonlinear evolution) remain non-negligible — only a factor of 2–3 smaller than projection effects. Our theoretical results agree fairly well with simulation measurements for the total bispectrum. To facilitate future applications and reproducibility, we make the corresponding code publicly available.
@@ 1 Introduction
The CMB bispectrum has been investigated with both WMAP (Fergusson et al., 2012) and Planck (Akrami et al., 2020a), though without significant PNG detection. In contrast, galaxy surveys have already yielded bispectrum measurements. The first detection came from the IRAS Point Source Catalog Redshift Survey (PSCz; Scoccimarro et al., 1998), with later surveys like BOSS (Gil-Mar´ın et al., 2015) refining the signal (Gil-Mar´ın et al., 2017; Cabass et al., 2022; D’Amico et al., 2022; Ivanov et al., 2023). However, these analyses have so far relied on the flat-sky approximation (all galaxy lines of sight are assumed to be parallel), the Limber approximation (see Limber, 1954; Kaiser, 1998; LoVerde and Afshordi, 2008) and Newtonian approximation where only the Newtonian gravitational dynamics (including RSDs) are considered, see Bernardeau et al. (2002). The next generation of galaxy surveys – especially those combining spectroscopic and photometric observations – will offer richer data and call for more sophisticated bispectrum analyses.
@@
In spectroscopic surveys such as the one of Euclid or DESI, precise redshift measurements with an error $\sigma_{z}=10^{-3}(1+z)$ (Mellier et al., 2025) enable the reconstruction of the three-dimensional galaxy distribution. Euclid, for instance, will collect spectroscopic data for $\sim$30 million galaxies over one-third of the sky (Mellier et al., 2025) – an order of magnitude more than BOSS and nearly the same number as that of DESI (Abdul Karim et al., 2025). Several frameworks have been developed to extract the bispectrum in redshift space, including Legendre polynomial decompositions like the Yamamoto–Scoccimarro formalism (Scoccimarro, 2015) and tripolar expansions (Sugiyama et al., 2019), supported by theoretical predictions (see e.g. Scoccimarro et al., 1999; Nan et al., 2018; de Weerd et al., 2020). However, as the survey volume increases, wide-angle effects become important, and the flat-sky approximation breaks down. These effects can be partially modeled within Fourier-space approaches (Castorina and White, 2018; Beutler et al., 2019; Castorina and di Dio, 2022; Noorikuhani and Scoccimarro, 2023; Pardede et al., 2023; Foglieni et al., 2023), but their inclusion in bispectrum analyses remains challenging. One solution to overcome the wide-angle effect is to use directly angular statistics. One well-suited approach for spectroscopic surveys is to use the spherical Fourier-Bessel (SFB) decomposition first introduced by Peebles (1973). Bertacca et al. (2018) derived the bispectrum expression, while Benabou et al. (2024) performed the numerical evaluation of the dominant terms. Additionally, on large scales, Newtonian approximation may become insufficient, motivating a check for relativistic effects as attempted in this work.
@@
Euclid will also carry out a wide-area photometric survey over the same sky region, detecting about 1.5 billion galaxies (Mellier et al., 2025). Due to limited redshift resolution ($\sigma_{z}=0.05(1+z)$), the data will be divided into a handful of tomographic bins. In such cases, the SFB decomposition is no longer optimal, as the poor radial resolution prevents a faithful recovery of the full 3D modes. Instead, these projected fields require angular statistics tailored to discrete redshift bins. This approach is sometimes referred as “tomographic spherical harmonics” (TSH), the approach that we consider in this work. Such angular statistics estimators are already well developed in the context of the CMB (see e.g. Akrami et al., 2020a) and can be adapted to galaxy clustering. For instance, Montandon et al. (2023) used the binned bispectrum estimator of Bucher et al. (2016). However, the theoretical prediction of the angular statistics of the number count involves slowly converging integrals over spherical Bessel functions. Mathematical techniques and efficient numerical methods for evaluating such integrals (e.g. efficient evaluation of hyper-geometric functions, FFTLog decompositions and suitable integration by part) has been studied in great details and are being applied to the power spectrum by Assassi et al. (2017) and Simonović et al. (2018) (see also Grasshorn Gebhardt and Jeong, 2018; Fang et al., 2020). The focus of the present work is to use the framework of Assassi et al. (2017) to compute the angular bispectrum of the matter number count.
@@
  • we ignore integrated projection effects such as lensing and Integrated Sachs–Wolfe. The computation of these contributions requires a more involved formalism with additional line-of-sight integrals, which we leave for future work. To enable comparisons, we remove these integrated effects from the relativistic simulation results.
@@ 2.1 First order
The first term of Eq. (4) represents the standard Newtonian perturbation theory, i.e. density plus redshift-space distortions, see Eq. (5). The remaining terms are due to relativistic effects. These effects are split into two different terms, one accounting for projection effects defined in Eq. (6) and another one for dynamical GR effects defined in Eq. (7). Let us stress that Eq. (4) is derived only under the assumption that photons travel along light-like geodesics and that matter motion is described by the Euler equation (i.e. galaxies follow time-like geodesics). Among the linear projection effects we have a Doppler term $\partial_{r}v_{1}$ which in the weak-field approximation is suppressed by a factor $\mathcal{H}/k$ with respect to density fluctuations, and terms which are directly proportional to the metric perturbations (as Sachs–Wolfe terms) which are suppressed by a factor $\left(\mathcal{H}/k\right)^{2}$ with respect to Newtonian perturbation theory.
@@ 2.2 Second order
For convenience, we split the whole expression
@@
where $\partial_{\alpha}$ is the angular derivative over the angle $\alpha=\{\theta,\varphi\}$ and we adopt Einstein’s index summation convention, into different pieces that have different physical origins such that
@@
(all terms are discussed in the remainder of this section), where we have defined the Newtonian number count perturbation at order $\mathcal{H}^{0}/k^{0}$, reading
@@ 3 The angular bispectrum
$C_{\ell}^{(n,m)}(r_{1})$ $=-\frac{2}{\pi}D\left(r_{1}\right){\mathcal{N}^{2}}\int drdk\ \tilde{W}_{r}D_{r}k^{4+n}P_{\phi_{0}}\left(k\right)j^{(m)}_{\ell}\left(kr\right)$
@@ 3.1 Pure second-order terms
The additional factor of $k_{1}^{-2}$ in the velocity (see Eq. (33)) shifts the powers of $k_{1}$. Consequently, one of the Newtonian terms (third line) now involves an integral over $r_{1}$ that does not simplify further. Moreover, taking into account that $f_{nm}^{(-4)}=0$ (see Eqs. B in Appendix B), this leaves only one pure GR contribution (first line) and a radiation contribution, analogous to Eq. (3.1), in the second line.
@@ 3.2 Quadratic terms
$\mathcal{A}_{\ell_{1}\ell_{2}\ell_{3}}=\frac{1}{2}\frac{\begin{pmatrix}\ell_{1}&\ell_{2}&\ell_{3}\\ 0&1&-1\end{pmatrix}+\begin{pmatrix}\ell_{1}&\ell_{2}&\ell_{3}\\ 0&-1&1\end{pmatrix}}{\begin{pmatrix}\ell_{1}&\ell_{2}&\ell_{3}\\ 0&0&0\end{pmatrix}}\,$. (43)
@@ 4 Analytical results
In Fig. 1, we plot each term of Eq. (44) for two redshift bins: $1.75\leq z\leq 2.25$ in the first row and $0.55\leq z\leq 0.65$ in the second row. The smoothness parameter $\sigma$ defined in Eq. (23) is chosen as a compromise: it must be small enough to satisfy the condition $\sigma\ll r_{\rm max}-r_{\rm min}$, ensuring a close match to a top-hat window, but large enough for the analytical derivatives to be numerically stable and well sampled. We find a good compromise with the criterion $r_{\rm max}-r_{\rm min}]/\sigma\simeq 0.05$. For $1.75\leq z\leq 2.25$ we set $\sigma=25$ Mpc/h/h while for $0.55\leq z\leq 0.65$ we set $\sigma=10$ Mpc/h/h. We indicate positive (negative) values with solid (dashed) lines. In black (dotted grey) we show the total GR (Newtonian) bispectrum. In the left column, we have fixed one multipole to a large value $\ell_{1}=512$ and vary $\ell_{2}=\ell_{3}$ so that the first point corresponds to a folded configuration with $2\ell_{2}=2\ell_{3}=\ell_{1}$ while the last point is an equilateral configuration that corresponds to the starting configuration of the second column. The middle column shows the equilateral configurations from $\ell_{1}=514$ to $\ell_{1}=2$ (note the flipped horizontal axis, with $\ell_{1}$ decreasing towards the right). Hence, the last point of the middle column corresponds to a large-scale triangle connecting with the first point of the third column where we fix $\ell_{1}=2$ and vary $\ell_{2}=\ell_{3}$. Therefore, the right part of the last column corresponds to squeezed triangles $\ell_{2}=\ell_{3}\gg\ell_{1}$.
@@ 5 Comparison with simulation
Since relativistic effects play only a subdominant role compared to Newtonian contributions, the total bispectrum can be compared with both Newtonian and relativistic simulations. For this we choose to compare the sum of all terms in Eq. (3), excluding the last three potential terms, with the bispectrum measured in the relativistic simulations. Regarding the relativistic effects, Montandon et al. (2023) estimated the “pure” relativistic contribution, including radiation, by subtracting the bispectrum measured in relativistic simulations from that measured in Newtonian simulations. However, this subtraction also incorporates several numerical artefacts that arise from the standard procedures used to include radiation and relativistic projection effects in Newtonian simulations. Specifically, two additional contributions must be considered in such a measurement:
  • Residual gauge effects: The relativistic ray tracer interprets particle positions in the Poisson gauge, even though in Newtonian simulations they are formally defined in the so-called $N$-body gauge Fidler et al. (2015). This mismatch has been estimated for the power spectrum by Montandon et al. (2023), where it was found to be subdominant relative to nonlinear relativistic corrections. Its impact on the bispectrum, however, has not been assessed before.
  • Residual backscaling effects: In Newtonian simulations, radiation is typically included by modifying the initial conditions so that the resulting power spectrum matches the correct radiation-included power spectrum at redshift $z=0$, while evolving only under matter gravitational dynamics. For the power spectrum, this effect has been estimated by Brandbyge et al. (2017) and Adamek et al. (2017), but we are not aware of a similar study for the bispectrum.
Therefore, the “pure” relativistic bispectrum obtained by Montandon et al. (2023) should be understood as quantifying the discrepancy between standard Newtonian simulations (employing backscaling) and relativistic simulations including the evolution of radiation. This discrepancy includes both physical contributions – namely the GR and radiation terms, $B^{D_{2}^{\rm GR}}_{\ell_{1}\ell_{2}\ell_{3}}+B^{D_{2}^{\rm Rad}}_{\ell_{1}\ell_{2}\ell_{3}}$ defined in Eq. (44) – and numerical artefacts (residual gauge and backscaling effects). Such a comparison can then disentangle which part of the numerical estimate originates from physical effects.
@@
In Fig. 6, we show the binned total bispectrum (in black) which includes all considered terms as dissected in Eq. (3) except the three last potential terms. In colours, we only show the dominant Newtonian contributions. The grey curves labelled with $\rm Simulation$’ are obtained by measuring the bispectrum in the relativistic simulations and setting the potential to zero for the ray tracing algorithm. The error bars indicate the standard error of the mean over the ten paired simulations. We obtain a quite good agreement between the theory and the simulation except at small scales. This was already observed in the power spectrum analysis of Montandon et al. (2023), and is due to the finite resolution of the simulations. Note that this explains why the discrepancy is stronger at lower redshift since each fixed multipole corresponds to a smaller scale. For the highest redshift bin, we observe good agreement for all configurations except for the smallest scales probed, while for the redshift bin $0.55\leq z\leq 0.65$, the measurements start to lose power with respect to the theory above $\ell\sim 100$. In the first Fig. 8, we show the results after applying smoothing with $\sigma_{\rm bin}=2$. At high redshift we observe excellent agreement, while at low redshift a small offset appears, with the simulations showing slightly less power even on large scales. This offset may arise from resolution effects, but could also be due to the smoothing procedure, which we discuss in more detail below.
@@
In Fig. 7, we show all the relativistic contributions to the bispectrum, as well as the measurements (in grey) obtained by subtracting the bispectrum of the relativistic and Newtonian simulations. It is important to note that the bispectrum is smoothed with $\sigma_{\rm bin}=1$, unlike in Fig. 6. In Fig. 8, we can also see the same results but with $\sigma_{\rm bin}=2$. Before comparing, we can note that for the redshift bin $1.75\leq z\leq 2.25$, the smoothing of the bispectrum tends to suppress the contribution of the GR part in violet ($D^{\rm GR}_{2}$) because it oscillates around zero. Hence, for that redshift, almost all dynamical effects are due to radiation. Otherwise, we recover the same kind of result as in Fig. 1. At lower redshift, both dynamical effects are always positive, which avoids any suppression. The sum of the two effects is then only about a factor $2$ smaller than the relativistic projection effect.
The measurement – quantifying the discrepancy between standard Newtonian simulations and consistent relativistic simulations – should be compared with the grey line obtained by summing the general relativistic and radiation effects. Note that the red crosses indicate error bars that include 0 (no detection). For the redshift $1.75\leq z\leq 2.25$, most large-scale triangles yield no detection. In the squeezed limit, we have a fairly good detection of the bispectrum in the simulations, but the analytical computation is about $3$ times smaller. This difference decreases for the lowest redshift bin $0.55\leq z\leq 0.65$, where the theory seems to agree in the squeezed triangle plot and in the large-scale equilateral configuration. However, we can see that almost all measurements are still above the theory prediction, indicating a slight overestimation in the simulation. For the small-scale equilateral triangles, the trend is similar for both redshifts, the theory decays strongly and then increases again to peak in the folded limit.
As explained at the beginning of this section, the purely relativistic part of the bispectrum measured from simulations contains both dynamical relativistic effects – including radiation – and numerical errors arising from the inconsistent gauge treatment and backscaling in Newtonian simulations. At the linear level, Montandon et al. (2023) showed that the gauge error introduced by the ray-tracing algorithm in Newtonian simulations is negligible. Similarly, Ref. Adamek et al. (2017) found that the residual backscaling effect contributes at the $\sim 1\%$ level on the largest scales of the matter power spectrum at low redshifts. Based on this, Montandon et al. (2023) concluded that their measurements were likely dominated by dynamical relativistic effects. Actually, based on our computations, we find that they are of the same order of magnitude as the numerical errors. This indicates that in the bispectrum, numerical artefacts can be as significant as the dynamical relativistic contributions.

### v3

@@ Abstract
The bispectrum of galaxy number counts is a key probe of large-scale structure, offering insights into the initial conditions of the Universe, the nature of gravity, and cosmological parameters. In this work, we present the first full-sky computation of the angular bispectrum in second-order perturbation theory without invoking the Limber approximation, and formulated for finite redshift bins via window functions, i.e., using tomographic spherical harmonics. To our knowledge, even the Newtonian part within this setup is novel. Building on this, we also include, up to second order in perturbation theory, the dynamical general relativistic and radiation effects, together with the leading relativistic projection effects. For simplicity, we neglect tracer bias and line-of-sight integrated contributions, however note that in particular the former can be straightforwardly incorporated within our framework. We evaluate the bispectrum contributions for two redshift bins, $1.75\leq z\leq 2.25$ and $0.55\leq z\leq 0.65$, and compare our theoretical prediction against relativistic light-cone simulations, with line-of-sight integral effects removed so as to enable direct consistency checks. As expected, we find that the Newtonian contributions are typically one or more orders of magnitudes larger than the relativistic signal across the entire spectrum for both redshifts. At $z=2$, we find that projection and dynamical relativistic effects have comparable amplitudes on large scales; somewhat unexpectedly, however, radiation effects dominate the relativistic signal in the squeezed limit. At $z=0.6$, the expected hierarchy is recovered, though dynamical corrections (from nonlinear evolution) remain non-negligible — only a factor of 2–3 smaller than projection effects. Our theoretical results agree fairly well with simulation measurements for the total bispectrum. To facilitate future applications and reproducibility, we make the corresponding code publicly available.
RBI-ThPhys-2025-01
@@ 1 Introduction
The CMB bispectrum has been investigated with both WMAP (Fergusson et al., 2012) and Planck (Akrami and others, 2020a), though without significant PNG detection. In contrast, galaxy surveys have already yielded bispectrum measurements. The first detection came from the IRAS Point Source Catalog Redshift Survey (PSCz; Scoccimarro et al., 1998), with later surveys like BOSS (Gil-Marín et al., 2015) refining the signal (Gil-Marín et al., 2017; Cabass et al., 2022; D’Amico et al., 2025; Ivanov et al., 2023). However, these analyses have so far relied on the local flat-sky approximation (Scoccimarro, 2015) — where the line of sight varies throughout the survey but each pair or triplet of galaxies is assumed to have parallel lines of sight when computing correlations — the Limber approximation (see Limber, 1954; Kaiser, 1998; LoVerde and Afshordi, 2008) and Newtonian approximation where only the Newtonian gravitational dynamics (including RSDs) are considered, see Bernardeau et al. (2002). The next generation of galaxy surveys – especially those combining spectroscopic and photometric observations – will offer richer data and call for more sophisticated bispectrum analyses.
@@
In spectroscopic surveys such as the one of Euclid or DESI, precise redshift measurements with an error $\sigma_{z}=10^{-3}(1+z)$ (Mellier and others, 2025) enable the reconstruction of the three-dimensional galaxy distribution. Euclid, for instance, will collect spectroscopic data for $\sim$30 million galaxies over one-third of the sky (Mellier and others, 2025) – an order of magnitude more than BOSS and nearly the same number as that of DESI (Abdul Karim and others, 2025). Several frameworks have been developed to extract the bispectrum in redshift space, including Legendre polynomial decompositions like the Yamamoto–Scoccimarro formalism (Scoccimarro, 2015) and tripolar expansions (Sugiyama et al., 2019), supported by theoretical predictions (see e.g. Scoccimarro et al., 1999; Nan et al., 2018; de Weerd et al., 2020). However, as the survey volume increases, wide-angle effects become important, and the flat-sky approximation breaks down. These effects can be partially modeled within Fourier-space approaches (Castorina and White, 2018; Beutler et al., 2019; Castorina and di Dio, 2022; Noorikuhani and Scoccimarro, 2023; Pardede et al., 2023; Foglieni et al., 2023), but their inclusion in bispectrum analyses remains challenging. Additionally, when correlating points at unequal redshifts, radial evolution effects stemming from the background should be taken into account (Bonvin et al., 2014; Beutler and Di Dio, 2020; Addis et al., 2025). One solution to overcome these effects within our framework is to use directly angular statistics integrated over redshift bins. Additionally, on large scales, Newtonian approximation may become insufficient, motivating a check for relativistic effects as attempted in this work.
@@
One well-suited approach — applicable to angular statistics — for spectroscopic surveys is to use the spherical Fourier-Bessel (SFB) decomposition first introduced by Peebles (1973). Bertacca et al. (2018) derived the bispectrum expression, while Benabou et al. (2024) performed the numerical evaluation of the dominant terms. Euclid will also carry out a wide-area photometric survey over the same sky region, detecting about 1.5 billion galaxies (Mellier and others, 2025). Due to limited redshift resolution ($\sigma_{z}=0.05(1+z)$), the data will be divided into a handful of tomographic bins. In such cases, the SFB decomposition is no longer optimal, as the poor radial resolution prevents a faithful recovery of the full 3D modes. Instead, these projected fields require angular statistics tailored to discrete redshift bins. This approach is sometimes referred as “tomographic spherical harmonics” (TSH), the approach that we consider in this work. Such angular statistics estimators are already well developed in the context of the CMB (see e.g. Akrami and others, 2020a) and can be adapted to galaxy clustering. For instance, Montandon et al. (2023) used the binned bispectrum estimator of Bucher et al. (2016). However, the theoretical prediction of the angular statistics of the number count involves slowly converging integrals over spherical Bessel functions. Mathematical techniques and efficient numerical methods for evaluating such integrals (e.g. efficient evaluation of hyper-geometric functions, FFTLog decompositions and suitable integration by part) has been studied in great details and are being applied to the power spectrum by Assassi et al. (2017) and Simonović et al. (2018) (see also Grasshorn Gebhardt and Jeong, 2018; Fang et al., 2020). The focus of the present work is to use the framework of Assassi et al. (2017) to compute the angular bispectrum of the matter number count.
@@
  • we ignore integrated projection effects such as lensing and Integrated Sachs–Wolfe. The computation of these contributions requires a more involved formalism with additional line-of-sight integrals, which we leave for future work. These terms can however become important. Di Dio et al. (2017) showed that lensing terms can become dominant for cross-correlations between well-separated redshift bins, while for auto-bispectra (all three redshifts equal), they can reach $10\%$ of the total amplitude for wide redshift bins. For simplicity, we therefore focus on the auto-bispectrum regime and enable comparisons by removing these integrated effects from the relativistic simulation results. Future work incorporating lensing terms would enable cross-bin-correlation studies and measurement of lensing signals through redshift separation techniques.
@@ 2.1 First order
The first term of Eq. (4) represents the standard Newtonian perturbation theory, i.e. density plus redshift-space distortions, see Eq. (5). The remaining terms are due to relativistic effects. These effects are split into two different terms, one accounting for projection effects defined in Eq. (6) and another one for dynamical GR effects defined in Eq. (7). Let us stress that Eq. (4) is derived only under the assumption that photons travel along light-like geodesics and that matter motion is described by the Euler equation (i.e. galaxies follow time-like geodesics). Among the linear projection effects we have a Doppler term $\partial_{r}v_{1}$ which in the weak-field approximation is suppressed by a factor $\mathcal{H}/k$ with respect to density fluctuations, and terms which are directly proportional to the metric perturbations (as Sachs–Wolfe terms) which are suppressed by a factor $\left(\mathcal{H}/k\right)^{2}$ with respect to Newtonian perturbation theory. These terms are included for completeness; however, since we consider projection effects only up to $\mathcal{H}/k$ at second order, we focus solely on the leading-order contributions in the following.
@@ 2.2 Second order
The second-order number count reads
@@
where $\partial_{\alpha}$ is the angular derivative over the angle $\alpha=\{\theta,\varphi\}$ and we adopt Einstein’s index summation convention. For convenience, we split the whole expression into different pieces that have different physical origins such that
@@
where all terms are discussed in the remainder of this section. The first term is the Newtonian number count perturbation at order $\mathcal{H}^{0}/k^{0}$, reading
@@ 3 The angular bispectrum
$C_{\ell}^{(n,m)}(r_{1})$ $=-\frac{2}{\pi}D\left(r_{1}\right){\mathcal{N}^{2}}\int drdk\ \tilde{W}_{r}D_{r}k^{4+n}P_{\phi_{0}}\left(k\right)j^{(m)}_{\ell}\left(kr_{1}\right)$
@@ 3.1 Pure second-order terms
The additional factor of $k_{1}^{-2}$ in the velocity (see Eq. 33) shifts the powers of $k_{1}$. Consequently, one of the Newtonian terms (third line) now involves an integral over $r_{1}$ that does not simplify further. Moreover, taking into account that $f_{nm}^{(-4)}=0$ (see Eqs. B in Appendix B), this leaves only one pure GR contribution (first line) and a radiation contribution, analogous to Eq. (3.1), in the second line.
@@ 3.2 Quadratic terms
$\mathcal{A}_{\ell_{1}\ell_{2}\ell_{3}}=\frac{1}{2}\left[\begin{pmatrix}\ell_{1}&\ell_{2}&\ell_{3}\\ 0&1&-1\end{pmatrix}+\begin{pmatrix}\ell_{1}&\ell_{2}&\ell_{3}\\ 0&-1&1\end{pmatrix}\right]\Bigg/\begin{pmatrix}\ell_{1}&\ell_{2}&\ell_{3}\\ 0&0&0\end{pmatrix}\,$. (43)
@@ 4 Analytical results
In Fig. 1, we plot each term of Eq. (44) for two redshift bins: $1.75\leq z\leq 2.25$ in the first row and $0.55\leq z\leq 0.65$ in the second row. The smoothness parameter $\sigma$ defined in Eq. (23) is chosen as a compromise: it must be small enough to satisfy the condition $\sigma\ll r_{\rm max}-r_{\rm min}$, ensuring a close match to a top-hat window, but large enough for the analytical derivatives to be numerically stable and well sampled. We find a good compromise with the criterion $\sigma/[r_{\rm max}-r_{\rm min}]\simeq 0.05$. For $1.75\leq z\leq 2.25$ we set $\sigma=25$ Mpc/h/h while for $0.55\leq z\leq 0.65$ we set $\sigma=10$ Mpc/h/h. We indicate positive (negative) values with solid (dashed) lines. In black (dotted grey) we show the total GR (Newtonian) bispectrum. In the left column, we have fixed one multipole to a large value $\ell_{1}=512$ and vary $\ell_{2}=\ell_{3}$ so that the first point corresponds to a folded configuration with $2\ell_{2}=2\ell_{3}=\ell_{1}$ while the last point is an equilateral configuration that corresponds to the starting configuration of the second column. The middle column shows the equilateral configurations from $\ell_{1}=514$ to $\ell_{1}=2$ (note the flipped horizontal axis, with $\ell_{1}$ decreasing towards the right). Hence, the last point of the middle column corresponds to a large-scale triangle connecting with the first point of the third column where we fix $\ell_{1}=2$ and vary $\ell_{2}=\ell_{3}$. Therefore, the right part of the last column corresponds to squeezed triangles $\ell_{2}=\ell_{3}\gg\ell_{1}$.
@@ 5 Comparison with simulation
Since relativistic effects play only a subdominant role compared to Newtonian contributions, the total bispectrum can be compared with both Newtonian and relativistic simulations. For this we choose to compare the sum of all terms in Eq. (3), excluding the last three potential terms, with the bispectrum measured in the relativistic simulations. Regarding the relativistic effects, Montandon et al. (2023) estimated the relativistic contribution by subtracting the bispectrum measured in relativistic simulations from that measured in Newtonian simulations. As we will discuss in more detail, these measurements cannot be reduced to a “pure” dynamical relativistic effects.
@@
In Fig. 6, we show the binned total bispectrum (in black) which includes all considered terms as dissected in Eq. (3) except the three last potential terms. In colours, we only show the dominant Newtonian contributions. The grey points labelled with $\rm Simulation$’ are obtained by measuring the bispectrum in the relativistic simulations and setting the potential to zero for the ray tracing algorithm. The error bars indicate the standard error of the mean over the ten paired simulations. We obtain a quite good agreement between the theory and the simulation except at small scales. This was already observed in the power spectrum analysis of Montandon et al. (2023), and is due to the finite resolution of the simulations. Note that this explains why the discrepancy is stronger at lower redshift since each fixed multipole corresponds to a smaller scale. For the highest redshift bin, we observe good agreement for all configurations except for the smallest scales probed, while for the redshift bin $0.55\leq z\leq 0.65$, the measurements start to lose power with respect to the theory above $\ell\sim 100$. In the first Fig. 8, we show the results after applying smoothing with $\sigma_{\rm bin}=2$. At high redshift we observe excellent agreement, while at low redshift a small offset appears, with the simulations showing slightly less power even on large scales. This offset may arise from resolution effects, but could also be due to the smoothing procedure, which we discuss in more detail below.
@@
In Fig. 7, we show all the relativistic contributions to the bispectrum, as well as the measurements (in grey dots). Note that the red plot markers come with error bars that include zero (which we do not show to not obscure the plot), indicating no detection. The bispectrum is smoothed with $\sigma_{\rm bin}=1$, unlike in Fig. 6. In Fig. 8, we can also see the same results but with $\sigma_{\rm bin}=2$. Before comparing, we can note that for the redshift bin $1.75\leq z\leq 2.25$, the smoothing of the bispectrum tends to suppress the contribution of the GR part in violet ($D^{\rm GR}_{2}$) because it oscillates around zero. Hence, for that redshift, almost all dynamical effects are due to radiation. Otherwise, we recover the same kind of result as in Fig. 1. At lower redshift, both dynamical effects are always positive, avoiding any suppression. The sum of the two effects is then only about a factor $2$ smaller than the relativistic projection effect.
The measurements are obtained by subtracting the bispectra of the relativistic and Newtonian simulations. For the redshift $1.75\leq z\leq 2.25$, most large-scale triangles yield no detection. In the squeezed limit, we have a significant detection of the bispectrum in the simulations. For the low redshift $0.55\leq z\leq 0.65$, almost all configurations are significantly detected.
The correct interpretation of these simulation measurements is tricky. Indeed, the subtraction incorporates several numerical artefacts that arise from the standard procedures used to include radiation and relativistic projection effects in Newtonian simulations. Specifically, two additional contributions must be considered in such a measurement:
  • Residual gauge effects: The relativistic ray tracer interprets particle positions in the Poisson gauge, even though in Newtonian simulations they are formally defined in the so-called $N$-body gauge (Fidler et al., 2015). This mismatch has been estimated for the power spectrum by Montandon et al. (2023), where it was found to be subdominant relative to nonlinear relativistic corrections. Its impact on the bispectrum, however, has not been assessed before.
  • Residual backscaling effects: In Newtonian simulations, radiation is typically included by modifying the initial conditions so that the resulting power spectrum matches the correct radiation-included power spectrum at redshift $z=0$, while evolving only under matter gravitational dynamics. For the power spectrum, this effect has been estimated by Brandbyge et al. (2017) and Adamek et al. (2017), and it was found to be $\sim 1\%$ on the largest scales of the matter power spectrum at low redshifts. We are not aware of a similar study for the bispectrum.
Therefore, the “pure” relativistic bispectrum obtained by Montandon et al. (2023) should be understood as the discrepancy between standard Newtonian simulations (employing backscaling) and relativistic simulations including the evolution of radiation. This discrepancy includes physical contributions — namely the GR and radiation terms, $B^{D_{2}^{\rm GR}}_{\ell_{1}\ell_{2}\ell_{3}}+B^{D_{2}^{\rm Rad}}_{\ell_{1}\ell_{2}\ell_{3}}$ defined in Eq. (44) — and numerical artefacts (residual gauge and backscaling effects). Based on the negligible magnitude of numerical artefacts at the linear level, Montandon et al. (2023) concluded that their measurements were likely dominated by dynamical relativistic effects.
We are now able to compute independently the dynamical relativistic and radiation effects. Their sum is shown in grey line in Fig. 7. As we can see, it is smaller by a factor $\sim 5$ than what was measured in the simulations of Montandon et al. (2023). Hence, we find that numerical errors are comparable to the physical effects. This indicates that in the bispectrum, numerical artefacts can be as significant as the dynamical relativistic contributions. Note that the discrepancy decreases for the lowest redshift bin $0.55\leq z\leq 0.65$, where the theory agrees better in the squeezed triangle plot and in the large-scale equilateral configuration. This can be explained by the numerical artefacts, since backscaling and gauge effects should decrease with the redshift.